OpenStudy (anonymous):

**WILL MEDAL** Which equations represent the line shown in the graph? Choose all answers that are correct. A.Y-1=-4/3(X+1) B.4X+3Y=-1 C.3X-4Y=1 D.Y=-4/3X-1/3

OpenStudy (immanuelv):

the way to check your answer is by plugging in the points in the graph to the answer. if they solve it then thats the answer :)

OpenStudy (anonymous):

Yea alot of ppl tell me to do that but idk how to

OpenStudy (immanuelv):

see the point (-2, 3)? you plug in (x, y) to your answers and see if left side = right side

OpenStudy (immanuelv):

(2, -3)* sorry

OpenStudy (immanuelv):

if (2, -3) works then you have to try (-1, 1). if both works thats your answer :)

OpenStudy (anonymous):

Sorry I'm confused

OpenStudy (immanuelv):

do you see the point (2, -3)? in the graph you provided?

OpenStudy (anonymous):

Yea

OpenStudy (immanuelv):

that point is pretty much (x, y) do you understand so far?

OpenStudy (anonymous):

Yea so it turn to be (2x,-3y)?

OpenStudy (immanuelv):

close... you substitute x for 2 and y for -3

OpenStudy (anonymous):

Ok

OpenStudy (immanuelv):

then you plug those in to the function for example Y-1=-4/3(X+1) (-3)-1 = 4/3(2+1)

OpenStudy (immanuelv):

-4 = 4 so the first answer choice will not work

OpenStudy (anonymous):

Ok

OpenStudy (immanuelv):

so if you do that with 4X+3Y=-1.... then you'll find out that B is the answer :)

OpenStudy (anonymous):

Would it be b and d?

OpenStudy (immanuelv):

it wont be d

OpenStudy (anonymous):

Ok so then it would be b and c then cuz u said that is wrong right?

OpenStudy (immanuelv):

d works

OpenStudy (immanuelv):

you were right... its b and d :) im sorry because i didnt see the - infront of -4/3

OpenStudy (anonymous):

Oh ok thx(:

OpenStudy (immanuelv):

see... you got this!!